OpenStudy (anonymous):

Is the following function periodic if yes what is its period: f(x) = [e^(j*pi*x/16)]*[cos(pi*x/17)] where "j" is complex

OpenStudy (anonymous):

\[f(x)= [e ^{(j*\pi*x)/16}]*[\cos(\pi* x/17)] \]

OpenStudy (thomas5267):

The first term has a period of 32, the second one has a period of 34.

OpenStudy (anonymous):

what is its period as whole function